# Depends-Kotlin: A Cross-Language Kotlin Dependency Extractor

### v1

@@ Abstract.
Since Google introduced Kotlin as an official programming language for developing Android apps in 2017, Kotlin has gained widespread adoption in Android development. However, compared to Java, there is limited support for Kotlin code dependency analysis, which is the foundation to software analysis. To bridge this gap, we developed Depends-Kotlin to extract entities and their dependencies in Kotlin source code. Not only does Depends-Kotlin support extracting entities’ dependencies in Kotlin code, but it can also extract dependency relations between Kotlin and Java. The extraction of such cross-language dependencies can help developers understand the migration process from Java to Kotlin. Additionally, we used a Java project with confirmed dependencies as a benchmark and converted this project to two projects: Kotlin-only and a combination of Kotlin and Java. The dependencies in these two projects were then extracted using our tool. The consistency observed among dependency relations in all three projects confirms the accuracy of Depends-Kotlin. Furthermore, the performance of Depends-Kotlin was assessed using another three projects of varying sizes. The source code of Depends-Kotlin and the dataset used in this demo paper have been uploaded to https://github.com/XYZboom/depends-kotlin. We also provided a screencast presenting Depends-Kotlin https://youtu.be/daZuXOwn1Ls.
Kotlin, Code Dependency, Cross-Language
copyright: acmcopyrightjournalyear: 2024doi: 10.1145/xxxxxxx.xxxxxxxconference: FSE’24: the 32nd ACM Symposium on the Foundations of Software Engineering (FSE ’24) 2024; November 15–19, 2024; Porto de Galinhas, Brazilbooktitle: Proceedings of the 32nd ACM Symposium on the Foundations of Software Engineering (FSE ’24), November 15–19, 2024, Porto de Galinhas, Brazilisbn: 978-1-4503-9999-9/18/06
@@ 1. Introduction
The dependency relations among entities in the source code form the foundations of software architecture analysis, including architecture recovery, architecture anti-pattern detection, architecture quality evaluation, and more (Mo et al., 2019; Cui et al., 2019; Xiao et al., 2022). Since Google introduced Kotlin as an official programming language for developing Android apps in 2017, Kotlin has gained widespread adoption in Android development. According to recent empirical studies, a significant of Android apps have been continuously migrated from Java to Kotlin (Ardito et al., 2020; Mateus and Martinez, 2020). However, compared to dependency analysis tools that support Java, such tool support for Kotlin is limited. To our knowledge, many well-known dependency analysis tools that support the Java language, such as Structure 101 (str, [n. d.]), Understand (und, [n. d.]), and DV8 (dv8, [n. d.]), currently do not offer support for the Kotlin language.
Kotlin dependency resolution faces two challenges. First, Kotlin was introduced by Google as “concise, expressive, and designed to be type and null-safe”. It contains a lot of syntactic sugar to ensure these features, thereby increasing the difficulty of resolving entity types and dependencies. Consider the following Kotlin code: Line 1 declares a class named Bar with a property x of type Int (property in Kotlin is similar to field in Java). Line 4 declares a high-level function named calculate that takes a lambda expression as its parameter. As shown in this line, the lambda takes a Bar type (called receiver type in Kotlin), and returns an Int. Line 7 declares another class named Foo, also with a property x of type Int. Line 8 declares calculateInFoo, which is a member function of the Foo class. It calls the calculate function in Line 9. Since the lambda parameter takes type Bar, the add function invoked in the lambda accesses the x property of the current Bar instance, not the Foo instance. The use of lambdas with receiver types as input allows concise syntax but it increases the difficulty of dependency resolution. In order to resolve this dependency (calculateInFoo use Bar.x), we need to locate calculate’s parameters and further trace them down to the correct type and properties. This is different from Java dependency analysis. Furthermore, Kotlin has its own dependency types, such as delegate and extension, which are supported by our tool and will be discussed in more detail in Section 2.
@@
3
4fun calculate(param: Bar.() -> Int) {}
5/* Function ’calculate’ takes a lambda with a receiver type Bar as its parameter.*/
6
7class Foo(val x: Int) {
@@
12}
Second, Kotlin is designed to be fully interoperable with Java and can run on the JVM, which makes it easy for developers to continuously migrate Java code to Kotlin. The following example includes a Kotlin class BarKotlin and a Java class FooJava. Lines 4-5 demonstrate that the Java class FooJava accepts and interacts with the Kotlin class BarKotlin by invoking a getX() method of the Kotlin class. However, in the Kotlin code, there is no explicit getX() method, as the getter method is implicitly generated by JVM for the class property. If we analyze only static code, dependencies between these two languages cannot be recognized. Considering that a significant percentage of Android apps in migration involve both Java and Kotlin code (Ardito et al., 2020; Mateus and Martinez, 2020), a tool for resolving Kotlin dependencies should not only address dependency relations in Kotlin code but also handle dependencies between Java and Kotlin code.
@@
To tackle these two challenges, we propose Depends-Kotlin, a cross-language Kotlin dependency extractor. Depends-Kotlin is based on the Depends framework, an open-source project designed for code dependency analysis (dep, [n. d.]; Jin et al., 2019). We enhanced Depends to support the Kotlin language by performing multi-round inference for getting Kotlin’s specific types and dependency relations. Additionally, we addressed cross-dependency relations between Kotlin and Java by refactoring the architecture of the original Depends framework and filling in the implicit code of Kotlin properties, which is handled by JVM and not shown in the source code.
To validate Depends-Kotlin’s accuracy, we manually converted a sample project written in Java into two separate projects: one exclusively containing only Kotlin code and the other featuring a combination of Kotlin and Java code. Using confirmed dependencies of Java code as the ground truth, we show that Depends-Kotlin can capture entities and their dependency relations accurately in both Kotlin code and Kotlin-Java scenarios. Additionally, we assessed Depends-Kotlin’s performance by analyzing the dependencies of another three Java-Kotlin projects with varing sizes.
@@ 2.1. Framework
In order to process Kotlin’s unique syntax and resolve its dependency types, we modified the existing Expression Analyser and Relation Resolver modules in Depends-Core. Specifically, We analyzed Kotlin’s extension functions and properties in Expression Analyser. We also identified the scope of Kotlin’s functions by adding related code as the context in Expression Analyser. Furthermore, we adapted Relation Resolver to handle dependencies with built-in types and new unique dependency types in Kotlin. For example, as method members of built-in types in Java always return built-in types, the original Depends chose to focus on the dependencies of analyzed source files and ignored built-in types. However, due to Kotlin extension functions, built-in types can return any types. This can greatly impact dependencies in the analyzed source code. Consequently we modified the Relation Resolver module to resolve such extension type.
For Kotlin’s entity parsing, we adopted the same logic of Java-parser in the original Depends framework. We used Antlr v4 to generate an Abstract Syntax Tree (AST) parser from a Kotlin grammar, which is provided on the official Kotlin website. This module is named as Kotlin Processor as shown in Figure  1. Subsequently, the Kotlin EntityTree Builder module store parsed AST nodes into concrete entities, such as files, packages, types, expressions, functions, and properties. In this step, some expressions’ types can be resolved in advance, which can reduce subsequent workload. As previously mentioned, the parser module of a specific language can work as a plug-in to the Depends-Core framework, enabling the analysis of other languages in the future.
Figure 1. Plug-in architecture of Depends-Kotlin.
2.2. Entity and Relation Resolution
To achieve accuracy with a good performance, we followed the same logic of the original Depends framework and performed multi-round inference for getting entities and dependency relations. An entity is represented by a tuple (id, name, entityType, context) and a relation is represented by (sourceId, targetId, dependencyType, weight). For simple dependency relations, such as extend, function parameters, function return and delegate, such entities and dependency relations are stored temporarily when analyzing expressions. For dependency relations, such as member access and member function call on expressions of unknown types, we conducted type inference. The basic idea is to deduct types from what is known to what is unknown. Any operation on an expression with a known type, such as a member function call, will yield the next expression with a known type. This process is recursively repeated until the entire expression’s type inference is completed. At this point, all entities and their dependencies should have been resolved.
It is worth mentioning that an additional search is required to resolve Kotlin extension functions during this process, as the scope of the expression containing extension functions needs to be identified. While traversing the syntax tree, if a function has a receiver type, it is marked as an extension function and the extended type is the receiver type. After the type system processing is complete, extension functions can be located based on their marks during the traversal of the entity tree. The extended types and the extension relationship of the function can then be recorded.
Our tool framework also automatically generates implicit code for Kotlin, and it distinguishes Java and Kotlin entities by assigning them different labels. When resolving expressions involving Kotlin-Java interactions, such cross-language relations can be easily detected. Table 1 lists the dependency types currently supported by Depends-Kotlin. The first six dependency relations can be observed in both Java-only and Kotlin-only projects. The delegate and extension relations are exclusive to Kotlin projects. The last two, JDependsOnK and KDependsOnJ, can be observed in projects containing dependencies between Java and Kotlin code.
Table 1. Relation types supported by Depends-Kotlin.
@@ 2.2. Entity and Relation Resolution
Extend a class extends a parent class or an interface
Call a statement in a method invokes another method
@@
JDependsOnK Java code depends on Kotlin entities
KDependsOnJ Kotlin code depends on Java entities
@@ 3. Evaluation
As Depends-Kotlin is designed for Kotlin and Kotlin-Java dependency resolution, we conducted an evaluation to assess its capability in resolving dependency relations in both Kotlin and Kotlin-Java scenarios. The evaluation focused on two perspectives: accuracy and performance.
@@ 3.1. Accuracy Verification
We selected a Java project with confirmed dependencies as the ground truth. Subsequently, we converted the Java project to Kotlin-only and Kotlin-Java projects. We utilized Depends-Kotlin to extract dependencies in these (partially) converted projects. As these three projects provide the same features, they are supposed to have similar entities and dependencies. If the extracted dependencies in the converted projects align with the ground truth, it proves that Depends-Kotlin can accurately capture Kotlin and Kotlin-Java dependency relations. If not, we should be able to explain the discrepancies, which may be due to the syntax difference between Java and Kotlin. The Java project we chose is implementation of a maze game (Cai et al., 2013) and in this game a maze comprises multiple rooms, each containing walls and a door leading to another room. The initial source code, named as ‘maze-java’, is written in Java and consists of 21 classes. The dependencies of ‘maze-java’ are extracted using the original Depends tool and have been confirmed as the ground truth.
We first converted all Java source code in ‘maze-java’ to Kotlin source code, named ‘maze-kotlin’. This process was automatically conducted by the conversion tool in IntelliJ IDEA. Next, we converted the source code of certain Java classes (e.g., data, utility) from ‘maze-java’ to Kotlin, resulting in ‘maze-kotlin-java’. The rationale behind choosing data/utility classes is to simulate real-world Java to Kotlin migration processes within a project  (Mateus et al., 2023). In this process, we converted 4 classes: Rectangle, Door, GreenRoom and RedRoom.
After the conversion, we applied Depends-Kotlin to analyze entities and entity relations of ‘maze-kotlin’ and ‘maze-kotlin-java’. We compared the results with those of ‘maze-java’, which is shown in Table 2. The upper part of this table shows the extracted entities at various granularities: Package, File, Class/Interface, Method Member, Java Field and Kotlin Property. The lower part displays the extracted dependency relations at the file level.
Table 2. Entity and relation extracted in maze game.
Entity maze-java maze-kotlin maze-kotlin-java
Package 2 2 2
File 21 21 21
Class, Interface 23 23 23
Method Member 73 17 69
Java Field 23 - 17
Kotlin Property - 31 6
Dependency Relation
Import 15 8 15
Extend 11 11 11
Call 176 121 163
Use 129 155 124
Parameter 48 41 48
Return 15 14 15
JDependsOnK - - 33
KDependsOnJ - - 9
For entities, the number of Package, File and Class is consistent across ‘maze-java’, ‘maze-kotlin’ and ‘maze-kotlin-java’. This alignment is expected as the general structure remains unchanged during the conversion. The number of Method Member decreased by 4, from 73 in ‘maze-java’ to 69 in ‘maze-kotlin-java’. It further decreased to 17 in ‘maze-kotlin’. The decrease of Method Member is because the getter/setter methods are converted into Kotlin property. This table also shows when 6 Java Field in ‘maze-java’ is converted, we have 6 Kotlin Property in ‘maze-kotlin-java’. In general, the number of Java fields should correspond to the number of converted Kotlin properties. However, the conversion algorithm prefers a property over a function when the method of Java is a getter and cheap to calculate [^1] https://kotlinlang.org/docs/coding-conventions.html#functions-vs-properties. In ‘maze-java’, an abstract class MapSite and its subclasses have a method named getColor, which is converted to as a read-only Kotlin property. Therefore, after the conversion, 8 additional Kotlin properties are generated in ‘maze-kotlin’.
For dependency relations, we observed that the number of extend relation has remained consistent among these three projects. For fully conversion from ‘maze-java’ to ‘maze-kotlin’, the number of import relation decreased from 15 to 8. This reduction is attributed to the fact that the Kotlin compiler can automatically conduct type inference, so there is no need to import the specified type. Furthermore, the number of call relations decreased by 55, from 176 to 121. This decrease can be explained by the conversion process, where Java methods are treated as Kotlin properties in ‘maze-kotlin’. Consequently, in ‘maze-kotlin’, accessing these Kotlin properties is categorized to the use relation. This also accounts for the increase in the number of use relations in ‘maze-kotlin’. The decrease in the number of parameter relation is also related to Kotlin directly replacing Java’s getter and setter methods with properties. This means that explicit method calls for accessing parameters are reduced in Kotlin. The number of return relation has remained consistent during the conversion process. This simplifies the process and allows a smooth transition to the converted code.
Our tool also detected the interactions between Java and Kotlin code: 33 JDependsOnK and 9 KDependsOnJ in ‘maze-kotlin-java’ as shown in Table 2. We manually inspected the code and found that Kotlin class extends Java class 3 times: Kotlin class Door extends Java class Mapsite, Kotlin class GreenRoom and RedRoom extend Java class Room. Besides, Kotlin class Door takes Java class Room as parameter 3 times, directly use it twice, and return it once. This finding confirms the accuracy of Depends-Kotlin in capturing Java-Kotlin interactions.
In summary, Depend-Kotlin demonstrates promise in the accuracy of entity and relation resolution in Kotlin and Kotlin-Java. The results indicate that the converted Kotlin entities can be accurately captured, and most of dependency relations in Kotlin code align with the relations in Java code. Furthermore, our findings suggest that, due to the assumption of the IDEA automatic migration process, dependency relations in Kotlin or Kotlin-Java code can differ from Java code.

### v3

@@ Abstract.
Since Google introduced Kotlin as an official programming language for developing Android apps in 2017, Kotlin has gained widespread adoption in Android development. However, compared to Java, there is limited support for Kotlin code dependency analysis, which is the foundation to software analysis. To bridge this gap, we develop Depends-Kotlin to extract entities and their dependencies in Kotlin source code. Not only does Depends-Kotlin support extracting entities’ dependencies in Kotlin code, but it can also extract dependency relations between Kotlin and Java. Using three open-source Kotlin-Java mixing projects as our subjects, Depends-Kotlin demonstrates high accuracy and performance in resolving Kotlin-Kotlin and Kotlin-Java dependencies relations. The source code of Depends-Kotlin and the dataset used have been made available at https://github.com/XYZboom/depends-kotlin. We also provide a screencast presenting Depends-Kotlin at https://youtu.be/ZPq8SRhgXzM.
copyright: acmcopyrightjournalyear: 2024doi: 10.1145/xxxxxxx.xxxxxxxconference: ASE’24: the 39th IEEE/ACM International Conference on Automated Software Engineering (ASE ’24); October 27 – November 01, 2024; Sacramento, California, USAbooktitle: Proceedings of the 39th IEEE/ACM International Conference on Automated Software Engineering (ASE ’24), October 27 – November 01, 2024, Sacramento, California, USAisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
The dependency relations among entities in the source code form the foundations of software architecture analysis (Mo et al., 2019; Cui et al., 2019; Xiao et al., 2022). Since Google introduced Kotlin as an official programming language for developing Android apps in 2017, Kotlin has gained widespread adoption in Android development. According to recent empirical studies, a large number of Android apps have been continuously migrated from Java to Kotlin (Ardito et al., 2020; Mateus and Martinez, 2020). However, compared to dependency analysis tools that support Java, such tool support for Kotlin is limited. To our knowledge, many well-known dependency analysis tools that support the Java language, such as Understand (und, [n. d.]), and DV8 (dv8, [n. d.]), currently do not offer support for the Kotlin language.
Kotlin dependency resolution faces two challenges. First, Kotlin is known as “concise, expressive, and designed to be type and null-safe”. It contains a lot of syntactic sugar to ensure these features, thereby increasing the difficulty of resolving entity types and dependencies. Consider the example in Listing 1: Line 1 declares a class named Bar with a property x of type Int (property in Kotlin is similar to field in Java). Line 3 declares a high-level function named calculate that takes a lambda expression as its parameter. As shown in this line, the lambda takes in a Bar type (called receiver type in Kotlin), and returns an Int. Line 5 declares another class named Foo, also with a property x of type Int. Line 6 declares calculateInFoo, which is a member function of the Foo class. It invokes the calculate function in Line 7. Since the lambda parameter takes type Bar, the add function invoked in the lambda accesses the x property of the current Bar instance, not the Foo instance. The use of lambdas with receiver types as input allows concise syntax but it increases the difficulty of dependency resolution. In order to resolve this dependency (calculateInFoo use Bar.x), we need to locate calculate’s parameters and further trace them down to the correct type and properties. Furthermore, Kotlin has its own dependency types, such as delegate and extension, which lacks support from existing tools..
@@
3fun calculate(param: Bar.() -> Int) {}
4/* Function ’calculate’ takes a lambda with a receiver type Bar as its parameter.*/
5class Foo(val x: Int) {
@@
10}
Listing 1: Kotlin syntax sugar example
Second, Kotlin is designed to be fully interoperable with Java and can run on the JVM, which makes it easy to continuously migrate Java code to Kotlin. Listing 2 includes a Kotlin class BarKotlin and a Java class FooJava. Lines 3-5 demonstrate that the Java class FooJava accepts and interacts with the Kotlin class BarKotlin by invoking a getX() method of the Kotlin class. However, in the Kotlin code the getter method is implicitly generated by JVM for the class property. If we analyze only static code, such dependencies cannot be recognized. Since a significant percentage of Android apps involve both Java and Kotlin code (Ardito et al., 2020; Mateus and Martinez, 2020), a tool for resolving Kotlin dependencies should not only address dependency relations in Kotlin but also handle dependencies between Java and Kotlin.
@@
Listing 2: Java-Kotlin implicit invocation
To tackle these two challenges, we propose Depends-Kotlin, a cross-language Kotlin dependency extractor. Depends-Kotlin is based on the Depends framework, an open-source project designed for code dependency analysis (Jin et al., 2019). We enhance Depends to support the Kotlin language by performing multi-round inference for getting Kotlin’s specific types and dependency relations. Additionally, we address cross-dependency relations between Kotlin and Java by refactoring the architecture of the original Depends framework and filling in the implicit code of Kotlin properties, which is handled by JVM and not shown in the source code. Depends-Kotlin is evaluated on three open-source projects containing both Kotlin and Java code and demonstrates its ability to accurately and efficiently extract Kotlin-Java and Kotlin-Kotlin dependencies.
Figure 1. Plug-in architecture of Depends-Kotlin
@@ 2.1. Framework
In order to process Kotlin’s unique syntax and resolve specific dependency relations, we modified the existing Expression Analyser and Relation Resolver modules in Depends-Core. Specifically, We analyzed Kotlin’s extension functions and properties in Expression Analyser. We also identified the scope of Kotlin’s functions by adding related code as the context in Expression Analyser. Furthermore, we adapted Relation Resolver to handle dependencies with built-in types and new unique dependency relations in Kotlin. For example, as method members of built-in types in Java always return built-in types, the original Depends chose to focus on the dependencies of analyzed source files and ignored built-in types. However, due to Kotlin extension functions, built-in types can return any types. This can greatly impact dependencies in the analyzed source code. Consequently we modified the Relation Resolver module to resolve such extension dependency relation.
For Kotlin’s entity parsing, we adopted the same logic of Java-parser in the original Depends framework. We used Antlr v4 to generate an Abstract Syntax Tree (AST) parser from a Kotlin grammar, which is provided on the official Kotlin website. This module is named as Kotlin Processor as shown in Figure 1. Subsequently, the Kotlin EntityTree Builder module store parsed AST nodes into concrete entities, such as files, packages, types, expressions, functions, and properties. In this step, some types in expressions can be resolved in advance, which can reduce subsequent workload. As previously mentioned, the parser module of a specific language can work as a plug-in to the Depends-Core framework, enabling the analysis of other languages in the future.
2.2. Entity and Dependency Relation Resolution
Table 1. Dependency Relations supported by Depends-Kotlin
@@ 2.2. Entity and Dependency Relation Resolution
Contain a class holds another class’s object as field
Extend a class extends a parent class
Implement a class implements an interface
Call an expression invokes another method
Create an expression in a method create an object
Cast an expression does a cast to a type
Annotation an entity uses an annotation
@@
Note: The first 11 dependency relations can be observed between Java and Kotlin entities. The last two are exclusive to Kotlin-Kotlin and Kotlin-Java.
Table 2. Dependency Relations Extracted by Depends-Kotlin
DSU-Sideloader Flap RootEncoder Total
K-J J-K K-K J-J K-J J-K K-K J-J K-J J-K K-K J-J
Import 19 0 170 1 3 0 170 1 102 190 913 193 1762
Contain 6 0 34 5 4 0 46 63 10 256 318 73 815
Extend 1 0 4 0 0 0 20 2 6 1 101 75 210
Implement 1 0 0 7 0 0 36 1 9 4 8 3 69
Call 50 0 450 81 3 0 370 479 62 423 1278 1316 4512
Create 1 0 31 2 3 0 94 25 59 71 489 57 832
Cast 0 0 0 1 5 0 10 15 0 8 48 9 96
Annotation 0 0 0 0 0 0 0 2 0 0 0 0 2
Use 70 0 703 187 9 0 713 1093 141 203 3331 4317 10767
Parameter 1 0 48 9 0 0 40 46 29 121 238 145 677
Return 3 0 36 16 0 0 71 24 2 30 263 30 475
Delegate 0 - 0 - 0 - 2 - 0 - 0 - 2
Extension 0 - 0 - 0 - 6 - 0 - 3 - 9
Total 152 0 1476 309 27 0 1578 1751 420 1307 6990 6218
Figure 2. J-K/K-J Dependency visualization(orange denotes K depends on J; green denotes J depends on K)
We followed the same logic of the original Depends framework and performed multi-round inference for getting entities and dependency relations. An entity is represented by a tuple (id, name, entityType, context) and a dependency relation is represented by (sourceId, targetId, dependencyType, weight). For simple dependency relations, such as extend, function parameter, function return and delegate, such dependency relations between two entities can be collected directly when analyzing expressions. For dependency relations, such as member access and member function call on expressions of unknown types, we conducted type inference. The basic idea is to deduct types from what is known to what is unknown. Any operation on an expression with a known type, such as a member function call, will yield the next expression with a known type. This process is recursively repeated until the entire expression’s type inference is completed.
An additional search is required to resolve Kotlin extension functions during this process, as the scope of the expression containing extension functions needs to be identified. While traversing the syntax tree, if a function has a receiver type, it is marked as an extension function and the extended type is the receiver type. After the type system processing is complete, extension functions can be located based on their marks during the traversal of the entity tree. The extended types and the extension relationship of the function can then be recorded.
Our tool framework also automatically generates implicit code for Kotlin, and it distinguishes Java and Kotlin entities by assigning them different labels. When resolving expressions involving Kotlin-Java interactions, such cross-language relations can be easily detected. Table 1 lists the dependency relations currently supported by Depends-Kotlin. Take Listing 2 for example, a Java expression is calling a Kotlin method getX() so a Call relation is extracted with Java source and Kotlin target information.
@@ 3. Evaluation
Subjects: Our subjects are DSU-Sideloader (1259 Stars, 94.1% Kotlin, 4.5% Java, 7k LOC), Flap (286 Stars, 52.4% Kotlin, 47.6% Java, 16k LOC) and RootEncoder (738 Stars, 55.8% Kotlin, 42.1% Java, 51k LOC). We chose these projects because they have a high number of stars and range in size from 7k to 51k LOC, allowing us to test our tool’s performance. Additionally, the Kotlin ratio varies among these three projects, from 94.1% in DSU-Sideloader to 55.8% in RootEncoder, which helps us test Depends-Kotlin’s ability to handle Kotlin-Kotlin and Kotlin-Java dependencies.
Run Depends-Kotlin: Following Figure 1’s architecture, Depends-Kotlin was implemented in Kotlin (the new Kotlin-Parser module) and Java (the legacy Depends-Core and Java-Parser modules). The README in the provided GitHub link specifies how to build and run this tool. Essentially, it takes the source file folder as input and outputs a JSON file with entities and dependency relations.
Results: Table 2 presents the extracted dependencies by Depends-Kotlin. The 1st column shows the dependency relations supported by our tool. The 2nd columns list the number of extracted dependencies in DSU-Sideloader, with each sub-column representing dependency relations from a particular source (Java or Kotlin) to a particular destination (Java or Kotlin). For example, 19 in the K-J column under DSU-Sideloader and the Import row means there are 19 instances where Kotlin files import a Java class, enum, or static method. 423 in the J-K column under RootEncoder and the Call row means there are 423 instances where an expression in Java code invokes a Kotlin method in RootEncoder.
Table 2 shows that all 13 dependency relations can be extracted from these three projects, with Use, Call, and Import being the top three most frequent. We also observed that Delegate and Extension have only 2 and 9 instances, respectively, indicating that these two new syntax features are not widely used in these three projects. The last row shows the total number of four kinds of relations in each project. There are 1476 K-K and 309 J-J relations in DSU-Sideloader with 94.1% Kotlin code, compared to 6990 K-K and 6218 J-J relations in RootEncoder with 55.8% Kotlin code. It makes sense that a higher ratio of Kotlin code tends to result in a higher ratio of K-K relations in a project. We also observed a high ratio of K-J (420) and J-K (1307) interactions in RootEncoder and significant K-J relation instances in DSU-Sideloader (152) and Flap (27), which demonstrates our tool’s ability to capture cross-language dependencies. Figure 2 presents the J-K (green arrow) and K-J (orange arrow) dependencies in the three subjects, with each node denoting a source file. As we can see, K-J and J-K dependencies are not limited to specific interfaces, demonstrating good interoperability between Java and Kotlin.
@@ 3.1. Accuracy Verification
To our knowledge, there is no available tool to directly extract Kotlin(-Java) dependency relations from source code. Although CodeQL (https://codeql.github.com) and DependencyPlugin (https://github.com/autonomousapps/dependency-analysis-gradle-plugin) can extract dependencies from Kotlin projects, they require compiling the entire project or only provide package-level dependencies, making them unsuitable for comparison with our tool.
Compiler Reference Check: We leveraged JetBrains’ Program Structure Interface (PSI) to conduct a first-round accuracy check. PSI is part of the Java/Kotlin compiler and can find references between elements in programs. If an extracted dependency instance by our tool can be found in references between two elements in PSI, we labeled it as “Found”; otherwise, we labeled it as “NotFound”. We calculate $Precision=\frac{Found\ Instances}{Found\ Instances\ +\ NotFound\ Instances}$ and $Recall=\frac{Found\ Instances}{Dependency\ Instances\ in\ PSI\ References}$ This is a coarse comparison, as PSI does not provide detailed dependency relations and we can only check the existence of extracted dependencies in PSI. Table 3 presents the precision value in our three subjects. Our tool shows promise in resolving dependencies both within the same language and across languages. Specifically, the average precision of J-J and K-K dependencies are 98.2% and 97.9%, respectively, while the average precision of K-J and J-K dependencies are 90.0% and 99.5%, respectively. Generally, the average precision of 97.9% indicates that the capture dependencies by Depends-Kotlin is precise. Due to space limit, the recall table is provided in the dataset link (see the abstract). The average recall across the three subjects is 85.0%, indicating that 85.0% of the dependencies in the ground truth can be correctly captured by Depends-Kotlin.
Manual Check: For each dependency type, we also randomly selected 5 instances from J-J, K-K, J-K, and K-J (if available) in our three subjects, and two authors with four years of Java/Kotlin experience conducted a thorough examination of the instances (the dataset is shared in the GitHub link provided in the abstract). They built the projects in IntelliJ IDEA and independently traced entities to check dependency instances. After completing this, they discussed the results until an agreement was reached. The manual inspection results served as ground truth and were compared with the dependency instances generated by Depends-Kotlin. The comparison shows that our tool can correctly capture 206 out of 213 dependencies (96.7% precision).
In general, both the compiler reference check and the manual check show that Depends-Kotlin can achieve good accuracy in capturing dependency relations.
Table 3. Precision Value of Three Projects
K-J J-K K-K J-J Average
Import 100.0% 100.0% 100.0% 100.0% 100.0%
Contain 100.0% 99.6% 99.4% 100.0% 99.6%
Extend 100.0% 100.0% 100.0% 100.0% 100.0%
Implement 100.0% 100.0% 100.0% 100.0% 100.0%
Call 89.5% 98.8% 98.1% 96.7% 97.4%
Create 100.0% 100.0% 98.2% 95.2% 98.2%
Cast 100.0% 100.0% 100.0% 100.0% 100.0%
Annotation - - - 100.0% 100.0%
Use 78.2% 99.5% 96.7% 98.7% 97.4%
Parameter 100.0% 100.0% 99.7% 99.5% 99.7%
Return 100.0% 100.0% 100.0% 85.7% 97.9%
Delegate - - 100.0% - 100.0%
Extension - - 88.9% - 88.9%
Average 90.0% 99.5% 97.9% 98.2% 97.9%